OpenStudy (anonymous):

Find the slope of the line passing through using the given points. (-3,2) and (1,-4)

OpenStudy (amistre64):

slope is found by subtracting one point from the other; it is the ration of y/x

OpenStudy (amistre64):

x y (-3 ,2 ) -( 1,-4 ) -------- -4 , 6 ; slope = y/x = -6/4 = -3/2